To make a 5.30 kg, 4.12 % (by mass) H2O2 solution, what is the amount (in g) of hydrogen peroxide (H2O2) needed?
Answer #1

% (w/w) H2O2 solution =

Given, %(w/w) = 4.12

Weight of solution = 5.30 kg = 5.301000 g

Then, = 4.12

Or, weight of H2O2 = (4.12*5.30*1000/100) = 218.36 g.

So, amount of Hydrogen peroxide needed = 218.36 g.